# NeuroLifting: Neural Inference on Markov Random Fields at Scale

### v2

@@ 4.4 Energy minimization with GNN
As indicated by Eq. 1, the energy function can serve as the loss function to guide network training since minimizing this energy function aligns with our primary objective. Typically, the energy function for a new problem instance takes the form of a look-up table, rendering the computation process non-differentiable. To facilitate effective training in a fully unsupervised setting, it is crucial to transform this computation into a differentiable loss aligning with the original energy function. The initial step involves transforming the decision variable from $x_{i}\in\{1,...,s_{i}\}$, where $s_{i}$ is the number of states of variable $x_{i}$, to $v_{i}\in\{0,1\}^{s_{i}}$. At any given time, exactly one element of the vector $v_{i}$ can be one, while all other elements must be zero; the position of the 1 indicates the current state of the variable $x_{i}$. Define $V_{k}=\otimes_{i\in C_{k}}v_{i}$, where $\otimes$ is the tensor product. The corresponding energy function would be Eq. 5. Subsequently, we relax the vector $v_{i}$ to $p_{i}(\theta)\in[0,1]^{s_{i}}$, where $p_{i}(\theta)$ represents the output of our network and $\theta$ denotes the network parameters. This output can be interpreted as the probabilities of each state that the variable $x_{i}$ might assume.
$E(\{v_{i}|i\in\mathcal{V}\})=\underbrace{\sum\limits_{i\in\mathcal{V}}\langle v_{i}(\theta),\phi(x_{i})\rangle}_{\text{Unary Term}}+\underbrace{\sum\limits_{C_{k}\in\mathcal{C}}\langle\psi(C_{K}),V_{k}\rangle}_{\text{Clique Term}}$
$L(\theta)=\underbrace{\sum\limits_{i\in\mathcal{V}}\langle p_{i}(\theta),\phi(x_{i})\rangle}_{\text{Unary Term}}+\underbrace{\sum\limits_{C_{k}\in\mathcal{C}}\langle\psi(C_{K}),P_{k}\rangle}_{\text{Clique Term}}$
@@
[…]
@@
Once the network outputs are available, we can easily determine the assignments by rounding the probabilities $p(\theta)$ to obtain binary vectors $v$. Using these rounded results, the actual energy can be calculated using Eq. 5. It is observed that after the network converges, the discrepancy between $L(\theta)$ and $E(\{v_{i}|i\in\mathcal{V}\})$ is minor and we won’t see any multi-assignment issue in decision variables. We choose Adam [55] as the optimizer, and employ simulated annealing during the training process, allowing for better exploring the loss landscape to prevent sub-optima.
@@ 4.5 Analysis and Discussion
Relation to lifting. In this innovative framework of using GNNs for inference on MRFs, a natural and sophisticated parallel emerges with the classical concept of lifting in optimization [20]. By mapping each unary term of an MRF to a node within a GNN and translating clique terms into densely connected subgraphs, the traditional MRF energy minimization transforms into optimizing a multi-layer GNN with extra dimensionality. This procedure aligns with the lifting technique where the problem space is expanded to facilitate more efficient computation. Akin to the principle of standard lifting to ease optimization, the GNN-based reparameterization can leverage the gradient descent optimization paradigm inherent in the smooth neural network landscape [56, 57], ensuring efficient computation and convergence. Therefore, while offering an enhanced approach to inference, the GNN reparameterization mirrors the core principles of lifting by transforming and extending the solution space into a computation-friendly one to achieve computational efficacy and scalability. More empirical evidence is in Section. 5.4.
Complexity analysis. The primary computations in this model arise from both the loss calculation and the operations within the GNN. For the loss function, let $c_{max}$ denote the maximum clique size. The time complexity for the loss calculation is given by $O(|\mathcal{V}||\mathcal{X}|+c_{max}|\mathcal{C}||\mathcal{X}|)$. For the GNN component, let $\mathcal{N}_{v}$ denote the average number of neighbors per node in the graph. The time complexity for neighbor aggregation in each layer is $O(\mathcal{N}_{v}|\mathcal{V}|)$ , and merging the results for all nodes requires $O(|\mathcal{V}|d)$ where $d$ is the feature dimension. Thus, for a $K$-layer GraphSAGE model with the custom loss function, the overall time complexity can be expressed as $O(|\mathcal{X}|(|\mathcal{V}|+c_{max}|\mathcal{C}|)+K|\mathcal{V}|(\mathcal{N}_{v}+d))$. This analysis highlights the efficiency of the framework in managing large-scale graphs by leveraging neighborhood sampling and aggregation techniques. The derived complexity indicates that the model scales linearly with respect to the number of nodes, the number of layers, and the dimensionality of the feature vectors, making it well-suited for large-scale instances.
@@ 5 Experiment
Baselines. We compare our approach against several well-established baselines: Loopy Belief Propagation (LBP), Tree-reweighted Belief Propagation (TRBP) [9], and Toulbar2 [12]. LBP is a widely used approximate inference algorithm that iteratively passes messages between nodes. TRBP improves upon LBP by introducing tree-based reweighing to achieve better approximations, particularly in complex graph structures. Toulbar2 is an exact optimization tool based on constraint programming and branch-and-bound methods Notably, Toulbar2 is the winner on all MPE and MMAP task categories of UAI 2022 Inference Competition  [^1] https://www.auai.org/uai2022/uai2022_competition. These baselines allow us to evaluate the performance of our proposed solution under fair settings.Note that comparisons with LBP and TRBP are omitted for high-order cases, as these methods are limited to simple scenarios on this kind of problems. We will use SRMP [35] on the high-order cases instead.
MRF format and transformation. The MRF data files are in UAI format and we interpret the data files in the same way as Toulbar2. Detailed information about unary and clique terms will be treated as unnormalized (joint) distributions, and the energies are calculated as $\theta_{i}(x_{i}=a)=-log(P(x_{i}=a))$, where $P(x_{i}=a)$ represents the probability provided by the data file. Note that we use the unnormalized values during the transformation process. The transformation for the clique energy terms will follow the same procedure. More details are in Appendix. I.
5.1 Synthetic Problems
We first conduct experiments on synthetic problems generated randomly based on Erdős–Rényi graphs [58]. The experiments are divided into pairwise cases and higher-order cases. We will compare the performance of NeuroLifting with LBP, TRBP, and Toulbar2 on pairwise MRFs. For the higher-order MRF cases, we will compare NeuroLifting exclusively with Toulbar2, as LBP and TRBP are not well-suited for handling the complexities inherent in high-order MRFs. The raw probabilities (energies) on the edges/cliques are randomly generated using the Potts function (Eq. 8), representing two typical types found in the UAI 2022 dataset. The parameters $\alpha$ and $\beta$ serve as constant penalty terms and $\mathbb{I}$ is the indicator function.
$\theta_{ij}=\alpha\mathbb{I}(x_{i}=x_{j})+\beta$ (8)
Pairwise instances. The inference results on pairwise cases are summarized in Table. 1. Due to the page limits here we only show the best results. The full table of the results with more statistics is shown in Table. 4 in Appendix. C. Prefix “P_potts_" and “P_random_" indicate instances generated with Potts energy and random energy, respectively. It is evident that as the problem size scales up, NeuroLifting outperforms the baseline approaches; meanwhile, it also achieves comparable solution quality even when the problem sizes are small. This trend is consistent across both energy models.
@@ 5.1 Synthetic Problems
Higher-order instances. The inference results on high order cases are summarized in Table. 2. The “H" in the prefix stands for High-order and all the instances are generated using Potts model. The number of cliques in the table encompasses both the cliques themselves and the edges connecting them. The relationships between nodes are based on either pairwise interactions or clique relationships. The results indicate that NeuroLifting outperforms Toulbar2 and SRMP, demonstrating its ability to effectively handle complicate high-order MRFs. This performance highlights the robustness and effectiveness of NeuroLifting across different graph structures.
Table 2: Results on the synthetic high order MRFs. Numbers correspond to the energy values. Best in bold. “NA" denotes that no solution was found within the specified time limits. Best in bold.
@@ 5.2 UAI 2022 Inference Competition datasets
We then evaluate our algorithm using instances from the UAI 2022 Inference Competition datasets, including both pairwise cases and high-order cases. The time settings will align with those established in the UAI 2022 Inference Competition, specifically 1200 seconds and 3600 seconds. LBP and TRBP algorithms are set to run for 30 iterations with a damping factor of 0.1, and the time limit for Toulbar2 is configured to 1200 seconds, which is generally sufficient for convergence. For NeuroLifting, we utilize an 8-layer GNN to model all instances, with the model trained for up to 100 iterations for each instance; other settings remain consistent with those used in the synthetic problems. We also experimented with lifting dimensions of 64, 512, 1024, 4096, and 8192.
Pairwise cases. We evaluate pairwise cases from the UAI MPE dataset. The results are displayed in Appendix. D. From Table. 5 in Appendix. D, we see that on trivial pairwise cases, where Toulbar2 successfully identifies the optimal solutions, NeuroLifting achieves comparably high-quality solutions that are on par with those obtained by LBP and TRBP. In cases where the problems become more challenging, although NeuroLifting does not surpass Toulbar2, it outperforms both LBP and TRBP. This suggests that NeuroLifting demonstrates improved performance on real-world datasets compared to simpler artificial instances. The full results of NeuroLifting with different lifting dimensions are detailed in Appendix. D.
High-order cases. For the high-order cases, we select a subset that has relatively large sizes. The results are presented in Table. 6 in Appendix. D. The performance of NeuroLifting aligns with the results obtained from synthetic instances, demonstrating superior efficacy on larger problems while consistently outperforming Toulbar2 in dense cases.
5.3 Physical Cell Identity
Physical Cell Identity (PCI) is an important parameter used in both LTE (Long-Term Evolution) and 5G (Fifth Generation) cellular networks. It is a unique identifier assigned to each cell within the network to differentiate and distinguish between neighboring cells. We transform PCI instances into pairwise MRFs, thus all the baselines could be evaluated. Appendix. J details how to perform the transformation.
We employ an internal real-world PCI data collection along with a synthetic PCI dataset for evaluation. The configurations for LBP, TRBP, and our proposed NeuroLifting approach are consistent with those outlined in Section. 5.1 expect for that we use 100 iterations here. For the Toulbar2 method, a time limit of 3600 seconds is set, while other parameters remain at their default values. The results are summarized in Table. 3. The first five instances are real-world PCI cases sourced from a city in China, while the latter five instances are generated. We see for smaller problem instances, Toulbar2 is able to solve them exactly. However, as the problem scale increases, it becomes increasingly challenging for Toulbar2 to effectively explore the solution space, and both LBP and TRBP struggle to converge. In contrast, NeuroLifting demonstrates strong generalization ability across all scales. Notably, it achieves commendable performance on large scales.
@@ 5.4 Analysis and Ablation Study
Efficiency Analysis Following UAI inference competition protocol, we compare NeuroLifting against Toulbar2 over a 1200-second horizon, collecting metrics at 200-second intervals. Due to Toulbar2’s logging limitations(no time information on complex problems), we evaluate it under varying time constraints. Results appear in Table 9 from Appendix G. Both methods perform efficiently on the first three instances, solving them within 200 seconds, with Toulbar2 showing stronger performance on these small-scale problems. However, as complexity increases, Toulbar2 fails to meet termination criteria within 1200 seconds and shows minimal solution quality improvement over time. In contrast, NeuroLifting maintains computational efficiency on larger instances and consistently outperforms Toulbar2 across all time intervals, delivering superior solutions.
Choice of GNN backbones. We evaluate model performance with various GNN backbones from Section 4.3, comparing their average results across UAI Inference Competition 2022 pairwise cases, our private PCI instances, and synthetic datasets (1000 nodes with average degree 4 or 8). Testing covers random energy configurations and Potts models. As shown in Fig. 5 in Appendix G, GraphSAGE consistently achieves superior results with fastest convergence across all datasets.
Choice of Optimizer.The optimizer selection is discussed in Section 4.4, based on problem structure analysis and empirical trials. We evaluate SGD, RMSprop, and Adam optimizers using UAI 2022 pairwise cases, with learning rate $10^{-4}$, embedded feature dimension 1024, and 8-layer network architecture. These configurations remain consistent across all test cases. Results with average loss curves in Fig. 6 from Appendix G demonstrate convergence differences, showing Adam outperforms both RMSprop and SGD in speed and stability.
Loss Landscape Visualization. We utilize the tool developed by [59] to visualize the loss landscape. Detailed settings of the visualization is in Appendix. H. We visualize the evolution of the loss landscape for networks with varying depths,specifically for $K\in\{1,2,5,8\}$. The resulting landscape visualizations are presented in Fig. 4 from Appendix. G, and the converged loss change trend in Fig. 7 from Appendix. G. We observe that a significant portion of the loss function is relatively flat, indicating that the loss can only decrease in constrained regions of the parameter space. As more layers are incorporated into the lifted model, it effectively expands these local regions, facilitating convergence toward better solutions. This characteristic suggests that the lifted model provides a greater capacity to navigate the optimization landscape.

### v3

@@ 4.4 Energy minimization with GNN
As indicated by Eq. 1, the energy function can serve as the loss function to guide network training since minimizing this energy function aligns with our primary objective. Typically, the energy function for a new problem instance takes the form of a look-up table, rendering the computation process non-differentiable. To facilitate effective training in a fully unsupervised setting, it is crucial to transform this computation into a differentiable loss aligning with the original energy function. The initial step involves transforming the decision variable from $x_{i}\in\{1,...,s_{i}\}$, where $s_{i}$ is the number of states of variable $x_{i}$, to $v_{i}\in\{0,1\}^{s_{i}}$. At any given time, exactly one element of the vector $v_{i}$ can be one, while all other elements must be zero; the position of the 1 indicates the current state of the variable $x_{i}$. Define $V_{k}=\otimes_{i\in C_{k}}v_{i}$, where $\otimes$ is the tensor product. The corresponding energy function would be Eq. 5. Subsequently, we relax the vector $v_{i}$ to $p_{i}(w)\in[0,1]^{s_{i}}$, where $p_{i}(w)$ represents the output of our network and $w\in W$ denotes the network parameters. This output can be interpreted as the probabilities of each state that the variable $x_{i}$ might assume.
$E(\{v_{i}|i\in\mathcal{V}\})=\underbrace{\sum\limits_{i\in\mathcal{V}}\langle v_{i},\phi(x_{i})\rangle}_{\text{Unary Term}}+\underbrace{\sum\limits_{C_{k}\in\mathcal{C}}\langle\psi(C_{K}),V_{k}\rangle}_{\text{Clique Term}}$
$L(W)=\underbrace{\sum\limits_{i\in\mathcal{V}}\langle p_{i}(w),\phi(x_{i})\rangle}_{\text{Unary Term}}+\underbrace{\sum\limits_{C_{k}\in\mathcal{C}}\langle\psi(C_{K}),P_{k}\rangle}_{\text{Clique Term}}$
@@
where $\langle\cdot,\cdot\rangle$ refers to the tensor inner product. The applied loss function is defined in Eq. 6, here $P_{k}=\otimes_{i\in C_{k}}p_{i}$. The rationale behind our loss function closely resembles that of the cross-entropy loss function commonly used in supervised learning. Let $P_{i}$ represent the true distribution and $Q_{i}$ denote the predicted distribution of the node $i$. A lower value of cross-entropy Eq. 7 indicates greater similarity between these two distributions. However, our approach differs in that we are not seeking a predicted distribution that closely approximates the true distribution. Instead, for each variable, we aim to obtain a probability distribution that is highly concentrated, with the concentrated points corresponding to the states that minimize the overall energy.
@@
Once the network outputs are available, we can easily determine the assignments by rounding the probabilities $p(w)$ to obtain binary vectors $v$. Using these rounded results, the actual energy can be calculated using Eq. 5. It is observed that after the network converges, the discrepancy between $L(W)$ and $E(\{v_{i}|i\in\mathcal{V}\})$ is minor and we won’t see any multi-assignment issue in decision variables. We choose Adam (Kingma & Ba, 2015) as the optimizer, and employ simulated annealing during the training process, allowing for better exploring the loss landscape to prevent sub-optima.
4.5 Theoretical Justification
We hereby discuss the logic of our method from a theoretical perspective. To enable differentiable optimization for the discrete MRF problem, we first lift the search space from discrete configurations to a continuous domain.
Definition 4.2 (Probabilistic Simplex).
Let $\mathcal{P}=\Delta_{1}\times\Delta_{2}\times\cdots\times\Delta_{|V|}$ be the Cartesian product of all distributions of nodes, where $h_{i}\in\Delta_{i}$ is a probability vector.
Based on this continuous domain, we define our specific neural parameterization. Instead of optimizing the variables directly, we parameterize them through a Graph Neural Network, allowing us to define the relaxed objective function.
Definition 4.3 (Neural Relaxation of MRFs).
NeuroLifting relaxes discrete $x_{i}$ into a probabilistic simplex $h_{i}\in\Delta^{K-1}$, where $h_{i}$ is obtained via a GNN with parameter $W$: $H(W)=\operatorname{Softmax}(\operatorname{GNN}_{W}(\mathcal{G}))$. The relaxed energy is defined as:
$\mathcal{L}(W)=\hat{E}(H(W))=\sum_{i\in V}\langle h_{i},\theta_{i}\rangle+\sum_{C\in\mathcal{C}}\sum_{y_{C}\in\mathcal{X}^{|C|}}\left(\prod_{j\in C}h_{j,y_{j}}\right)\theta_{C}(y_{C})$
A fundamental requirement for any relaxation method is consistency: we must ensure that minimizing the new continuous objective yields the same solution as the original discrete problem, without introducing a “relaxation gap.”
{restatable}
propositioneqrelation(Consistency of Relaxation) The minimal relaxed energy Eq. 8 over continuous $\mathcal{P}$ is equal to the minimal energy Eq. 5 over the discrete domain $\mathcal{X}$, and is reached at some vertices of $\mathcal{P}$.
$\min_{H\in\mathcal{P}}\hat{E}(H)=\min_{x\in\mathcal{X}}E(x)$ (9)
While Proposition 4.3 guaranties that global optimality coincides, the relaxed energy landscape $\hat{E}(H)$ remains non-convex and may contain high energy barriers between local minima. This is where the advantage of neural parameterization becomes evident. We first assume the network is sufficiently expressive.
Assumption 4.4 (Structural Homogeneity and Scaling).
The neural architecture $\Phi(W)$ is locally homogeneous with respect to weight scaling. As the parameter norm approaches zero ($W||\to\epsilon$), the logits $Z$ converge to zero, forcing the Softmax output $H(W)$ to the uniform distribution $U$.
Assumption 4.5 (High-Dimensional Manifold Connectivity).
The lifting dimension satisfies $dim(\mathcal{W})\gg dim(\mathcal{P})$. This redundancy implies that for any configuration $H\in\mathcal{P}$, the fiber $\Phi^{-1}(H)\subset\mathcal{W}$ is a high-dimensional manifold, providing sufficient degrees of freedom to navigate the energy landscape.
Under these assumptions, we can demonstrate that the over-parameterized space $\mathcal{W}$ possesses superior topological properties compared to the variable space $\mathcal{P}$, specifically regarding the connectivity of solutions.
{restatable}
theoremconnectivity(Connectivity of Parameter Space) For any two discrete states $H^{(a)}$ and $H^{(b)}$, even though the linear interpolation between them may pass a high energy barrier, there exists a trajectory in the GNN parameter space $W(t)\in\mathcal{W},t\in[0,1]$ connecting $W_{a}$ and $W_{b}$ (where $H(W_{a})\approx H^{(a)}$ and $H(W_{b})\approx H^{(b)}$). Furthermore, one can “bypass” the energy barrier by controlling the norm of parameters.
Remark 4.6.
Note the constructive proof of Theorem 4.5 only exhibits the existence of the detour optimization trajectories. In practice, the optimization dynamics may follow different behavior. This theorem is consistent with the phenomenon of mode connectivity observed in deep learning (Garipov et al., 2018). Our neural parameterization lifts the optimization landscape, creating low-energy tunnels between solutions that are separated by high barriers in the discrete space.
Having established that the landscape admits barrier-free trajectories, we finally address the stability of the optimization process. The following theorem guarantees that gradient-based updating will settle to a stationary point.
{restatable}
theoremconvergence(Convergence of NeuroLifting) Define $\mathcal{L}(W)=\hat{E}(H(W))$ as the loss function. Assume $\mathcal{L}$ is $L_{total}$-Lipschitz continuous. If the learning rate satisfies $0<\eta\leq\frac{2}{L_{total}}$, then the optimization converges to a stationary point.
Proof of Proposition 4.3, Theorem 4.5 and Theorem 4.6 can found in Appendix B.1, B.2 and B.3, respectively.
@@ 4.6 Analysis and Discussion
Relation to lifting. Our framework aligns with the classical concept of lifting (Balas et al., 1993) by expanding the solution space to facilitate optimization. By reparameterizing MRF terms into a GNN, we transform the rugged discrete energy minimization into an optimization over a smooth, high-dimensional neural landscape (Dauphin et al., 2014; Choromanska et al., 2015). This “NeuroLifting” enables the use of efficient gradient descent algorithms, leveraging the favorable optimization properties of deep networks to achieve better scalability and convergence than traditional discrete methods.
Complexity analysis can be found in Appendix. C
@@ 5 Experiment
Baselines. We compare our approach against several well-established baselines: Loopy Belief Propagation (LBP), Tree-reweighted Belief Propagation (TRBP) (Wainwright et al., 2005), and Toulbar2 (De Givry, 2023). LBP is a widely used approximate inference algorithm that iteratively passes messages between nodes. TRBP improves upon LBP by introducing tree-based reweighing to achieve better approximations, particularly in complex graph structures. Toulbar2 is an exact optimization tool based on constraint programming and branch-and-bound methods Notably, Toulbar2 is the winner on all MPE and MMAP task categories of UAI 2022 Inference Competition  [^1] https://www.auai.org/uai2022/uai2022_competition. These baselines allow us to evaluate the performance of our proposed solution under fair settings.Note that comparisons with LBP and TRBP are omitted for high-order cases, as these methods are limited to simple scenarios on this kind of problems. We will use SRMP (Kolmogorov, 2015) on the high-order cases instead.
MRF format and transformation. The MRF data files are in UAI format and we interpret the data files in the same way as Toulbar2. Detailed information about unary and clique terms will be treated as unnormalized (joint) distributions, and the energies are calculated as $\theta_{i}(x_{i}=a)=-log(P(x_{i}=a))$, where $P(x_{i}=a)$ represents the probability provided by the data file. Note that we use the unnormalized values during the transformation process. The transformation for the clique energy terms will follow the same procedure. More details are in Appendix J.
@@
5.1 Synthetic Problems
We first conduct experiments on synthetic problems generated randomly based on Erdős–Rényi graphs (Erdös & Rényi, 1959). The experiments are divided into pairwise cases and higher-order cases. We will compare the performance of NeuroLifting with LBP, TRBP, and Toulbar2 on pairwise MRFs. For the higher-order MRF cases, we will compare NeuroLifting exclusively with Toulbar2, as LBP and TRBP are not well-suited for handling the complexities inherent in high-order MRFs. The raw probabilities (energies) on the edges/cliques are randomly generated using the Potts function (Eq. 10), representing two typical types found in the UAI 2022 dataset. The parameters $\alpha$ and $\beta$ serve as constant penalty terms and $\mathbb{I}$ is the indicator function.
$\theta_{ij}=\alpha\mathbb{I}(x_{i}=x_{j})+\beta$ (10)
Pairwise instances. The inference results on pairwise cases are summarized in Table. 1. Due to the page limits here we only show the best results. The full table of the results with more statistics is shown in Table 4 in Appendix D. Prefix “P_potts_” and “P_random_” indicate instances generated with Potts energy and random energy, respectively. It is evident that as the problem size scales up, NeuroLifting outperforms the baseline approaches; meanwhile, it also achieves comparable solution quality even when the problem sizes are small. This trend is consistent across both energy models.
Higher-order instances. The inference results on high order cases are summarized in Table. 2. The “H” in the prefix stands for High-order and all the instances are generated using Potts model. The number of cliques in the table encompasses both the cliques themselves and the edges connecting them. The relationships between nodes are based on either pairwise interactions or clique relationships. The results indicate that NeuroLifting outperforms Toulbar2 and SRMP, demonstrating its ability to effectively handle complicate high-order MRFs. This performance highlights the robustness and effectiveness of NeuroLifting across different graph structures.
Table 2: Results on the synthetic high order MRFs. Numbers correspond to the energy values. Best in bold. “NA” denotes that no solution was found within the specified time limits. Best in bold.
@@ 5.2 UAI 2022 Inference Competition datasets
We evaluate our algorithm on the UAI 2022 Inference Competition datasets, including both pairwise and high-order cases, with time limits of 1200 and 3600 seconds as specified in the competition. LBP and TRBP run for 30 iterations with a damping factor of 0.1, while Toulbar2 is limited to 1200 seconds. For NeuroLifting, we use an 8-layer GNN trained for up to 100 iterations per instance, with lifting dimensions of 64, 512, 1024, 4096, and 8192; other settings follow the synthetic experiments.
Pairwise cases. We evaluated pairwise cases from the UAI MPE dataset, with results in Appendix E. Table 5 shows NeuroLifting achieves solutions comparable to LBP and TRBP on trivial problems where Toulbar2 finds optimal solutions. On more challenging problems, while not surpassing Toulbar2, NeuroLifting outperforms both LBP and TRBP, indicating better performance on real-world datasets than artificial instances. Complete results with varying lifting dimensions appear in Appendix E.
High-order cases. For the high-order cases, we select a subset that has relatively large sizes. The results are presented in Table 6 in Appendix E. The performance of NeuroLifting aligns with the results obtained from synthetic instances, demonstrating superior efficacy on larger problems while consistently outperforming Toulbar2 in dense cases.
5.3 Real-world Challenge: Physical Cell Identity
Physical Cell Identity (PCI) uniquely identifies cells in LTE and 5G networks, distinguishing between neighboring cells. We transform PCI instances into pairwise MRFs to enable evaluation across all baselines. Transformation details are provided in Appendix K. We evaluated using internal real-world PCI data and synthetic datasets. LBP, TRBP, and NeuroLifting configurations match Section 5.1, but with 100 iterations, while Toulbar2 had a 3600-second limit with default parameters. Table 3 shows results for five real-world cases from a Chinese city and five synthetic instances. Toulbar2 solves smaller problems exactly but struggles with larger scales. Similarly, LBP and TRBP face convergence issues on complex problems. NeuroLifting demonstrates strong generalization across all scales, achieving notable performance even on large instances.
@@ 5.4 Analysis and Ablation Study
Efficiency Analysis. Following UAI protocol, we compared NeuroLifting against Toulbar2 over 1200 seconds, with metrics at 200-second intervals (due to Toulbar2’s logging limitations, there is no time information on complex problems). Results in Table 9 (Appendix H) show Toulbar2 performs better on simpler instances, solving the first three within 200 seconds. However, on complex problems, Toulbar2 fails to terminate within 1200 seconds with minimal quality improvement. NeuroLifting maintains efficiency on larger instances, consistently outperforming Toulbar2 across all time intervals with superior solutions.
Choice of GNN backbones. We evaluated GNN backbones from Section 4.3 across UAI 2022 pairwise cases, private PCI instances, and synthetic datasets (1000 nodes, average degree 4/8), testing both random energy configurations and Potts models. Fig. 5 in Appendix H demonstrates GraphSAGE’s consistent superiority in both results quality and convergence speed across all datasets.
Choice of Optimizer. Optimizer selection, discussed in Section 4.4, is based on problem structure analysis and empirical testing. We evaluated SGD, RMSprop, and Adam on UAI 2022 pairwise cases using learning rate $10^{-4}$, 1024-dimensional embedded features, and 8-layer networks across all tests. Results in Fig. 6 (Appendix H) show Adam’s superior convergence speed and stability compared to RMSprop and SGD.
Loss Landscape Visualization. We visualize loss landscapes using the tool from Li et al. (2018), with detailed settings in Appendix I. Fig. 4 in Appendix H shows landscape evolution for networks of depths $K\in\{1,2,5,8\}$, while Fig. 7 displays converged loss trends. We observe that much of the loss function remains flat, with decreases possible only in limited parameter space regions. Deeper lifted models effectively expand these regions, enabling better solution convergence and demonstrating enhanced optimization landscape navigation capacity.